OpenStudy (sophadof):

Quadratic functions anyone? h(x)=-x^2-6x-9 f(x)=2x^2+9x+4 g(x)=x^2+x-20

OpenStudy (wolf1728):

Do you just want those solved?

OpenStudy (anonymous):

I can solve it but its gonna take a lil min

OpenStudy (anonymous):

maybe lik 10 min

OpenStudy (anonymous):

first those equations have to be a x and y not both x...are you sure those are the right equations

OpenStudy (sophadof):

yes @lovingod786

OpenStudy (sophadof):

@lovingod786

OpenStudy (sophadof):

@wolf1728 Yeah..but I need to show a little work..

OpenStudy (anonymous):

Ik im doing it rn

OpenStudy (anonymous):

Ok so you're gonna have double arenthese like (x 3) (x 3)...all you have to do is figure out the signs like negative and positive to get the equation lik -x^2 -6x-9

OpenStudy (anonymous):

nvm...ima just do it the long way

OpenStudy (sophadof):

haha... I was confused for a second @lovingod786

OpenStudy (anonymous):

so first youre gonna identify your A,B, and C in the equation....ax^2+bx+c=0

OpenStudy (anonymous):

Then you're gonna plug those letters into the equation

OpenStudy (sophadof):

would -x just be -1? @lovingod786

OpenStudy (anonymous):

\[x=\frac{ -b+\sqrt{b^2 -4ac}}{ 2a }\]

OpenStudy (anonymous):

No it wudnt...ur solving for x

OpenStudy (anonymous):

so in the equation where it states -b..it would only b -b if your b from the original equation is positve.

OpenStudy (sophadof):

but the b from the original is negative...

OpenStudy (anonymous):

so therefore it wud b positve

OpenStudy (anonymous):

But the B that is in the sqrt root wud b the original negative B

OpenStudy (anonymous):

This is how your equation should look

OpenStudy (anonymous):

\[x= \frac{ 6 + \sqrt{(-6^2) - 4 (-1)(-9)} }{2(-1)}\]

OpenStudy (anonymous):

get it?

OpenStudy (sophadof):

YES!!(:

OpenStudy (anonymous):

You can finish it from here?

OpenStudy (sophadof):

I will...try?

OpenStudy (anonymous):

Oh and btw where the + is b4 he sqaure root...how I was taught was..its also a minus sign under the addition sign so you will come out with two answer and either or.

OpenStudy (anonymous):

Lemme see you work ou tthe problem

OpenStudy (anonymous):

The either or is basically the same thing but one you get the subtraction you solve for the addition with the same stuff

OpenStudy (anonymous):

\[\frac{ a }{ b } + \frac{ a }{ b }\]

OpenStudy (anonymous):

\[\frac{ a }{ b } - \frac{ a }{ b }\]

OpenStudy (anonymous):

Thats basically the either or format

OpenStudy (anonymous):

Do you have the answer??

OpenStudy (sophadof):

no.. I have no idea what to do now @lovingod786

OpenStudy (sophadof):

I thought I could but..

OpenStudy (anonymous):

What happened

OpenStudy (anonymous):

Where did you get stuck

OpenStudy (sophadof):

I just don't know what to do... I don't know where to start and what to start..

OpenStudy (anonymous):

Ok just look...ima gonna do step by step.

OpenStudy (sophadof):

ok! I will write this down so I can look later on too!

OpenStudy (anonymous):

\[x = \frac{ -b + \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

Wait...do you understand where the A,B, and C.come fromwith the original equation?

OpenStudy (sophadof):

yes!

OpenStudy (anonymous):

ok i will continue now

OpenStudy (anonymous):

\[x= \frac{ 6 + \sqrt{-6^2 - 4(9)} }{ -2 }\]

OpenStudy (anonymous):

\[x = \frac{ 6+ \sqrt{-6^2 - 36} }{ -2 }\]

OpenStudy (anonymous):

\[x = \frac{ 6 + \sqrt{36-36} }{ -2 }\]

OpenStudy (anonymous):

\[x = \frac{ 6 + 0 }{ -2 }\]

OpenStudy (anonymous):

\[x = \frac{ 6 }{ -2 } = -3\]

OpenStudy (sophadof):

THANK YOU! That helped a lot!

OpenStudy (sophadof):

When I do the other problems (tomorrow, it's late), would I be able to check my work with you? @lovingod786

OpenStudy (anonymous):

Yeah just tag me like you're doin now

OpenStudy (anonymous):

I will get a notification

OpenStudy (sophadof):

Will do! So is that the final answer, -3?

OpenStudy (anonymous):

Yeah

OpenStudy (sophadof):

Yay! That was so much easier than I expected... You were very good at explaining! thank you!

OpenStudy (anonymous):

You're welcome

OpenStudy (sophadof):

does x=15.75 for b? @lovingod786

OpenStudy (anonymous):

lol oh snap I didnt even plug inhold on

OpenStudy (sophadof):

hah! it's ok!(:

OpenStudy (anonymous):

\[h(x) = -x^2 -6x -9\]

OpenStudy (anonymous):

\[h(-3)=-(-3)^2 -6(-3) - 9\]

OpenStudy (anonymous):

\[h(-3) = -(9) + 18 + 9 \]

OpenStudy (anonymous):

Answer should b 0

OpenStudy (sophadof):

Dang it... so now what?

OpenStudy (sophadof):

I think I am going to go to sleep and try again in the morning!! for b: x=15.75 for c: I didn't understand @lovingod786 Good night!

OpenStudy (sophadof):

thank you anyways!

OpenStudy (anonymous):

I got the wrong answer hold on

OpenStudy (anonymous):

You should follow PEDMAS

OpenStudy (anonymous):

So it should be..

OpenStudy (anonymous):

\[h(-3) = (3)^2 +18 -9\]

OpenStudy (anonymous):

then so exponents so the equation wud b 9+9 = 18

OpenStudy (anonymous):

I even checked with mathway

OpenStudy (anonymous):

18 is the final answer for the equation h(18)

OpenStudy (wolf1728):

h(x)=-x^2-6x-9 testing the -3 answer h(x) = -(-3^2) -6*-3 -9 h(x) = -9 +18 -9 h(x) =0 so -3 is a valid answer

OpenStudy (wolf1728):

lovingod786 you used the quadratic formula for equation (h) x = [6 +-sq root (36 -36)] / -2 x = 6 +- sq root(0) / -2 -3 is one valid root what is the other root?

OpenStudy (anonymous):

What do you mean the other root? @wolf1728

OpenStudy (anonymous):

Because since it is a zero you dont need two answers

OpenStudy (wolf1728):

Solving the equation with the quadratic formula: x = 6 +- sq root(0) / -2 Notice there is a "+-" before the sq root.

OpenStudy (wolf1728):

Yes, so both answers are -3 and -3

OpenStudy (anonymous):

There's no two answers for this...because all you are doing is subtracting and adding the zero so you could just leave it as -3 but if you want to write it like how you doing it then go right ahead

OpenStudy (wolf1728):

The only reason I am stating both answers is that teachers always insist upon the student giving BOTH answers (even if the answers are equal).

OpenStudy (anonymous):

My bad...You don't plug intothe equation as I see...The answer you got where the formula contains a square root is the answer

OpenStudy (wolf1728):

It seems you were testing the -3 answer with the equation h(−3)=(3)2+18−9 and then you say 18 is the final answer for the equation h(18) Actually, you miscalculated in setting up h(−3)=(3)2+18−9 it should be h(−3)=-(-3)2 -(6*-3) −9 h(−3)= -9 +18 -9 h(-3) = 0

OpenStudy (wolf1728):

Yes, plugging the number into the equation is just the way to test if the answer is correct

OpenStudy (anonymous):

I got that answer at first but then I plugged it into mathway to double check and they gave me 18 for the whole equation

OpenStudy (wolf1728):

Well at least we have the answer for the first equation :-)

OpenStudy (anonymous):

Lol need help with the rest??

OpenStudy (wolf1728):

I think we should stop here and maybe have sophadof do a little work too.

OpenStudy (anonymous):

ok

OpenStudy (wolf1728):

ok see ya

OpenStudy (sophadof):

I am working on it!!!(: sorry for not responding, i was at work @wolf1728 @lovingod786

OpenStudy (wolf1728):

Here's the second quadratic equation f(x)=2x^2+9x+4 with a, b and c put into the quadratic formula: x= [-9 +-sq root(81 -4*2*4)] / 2*2 Can you solve it from here?